# Interactive Agents to Overcome Ambiguity in Software Engineering

### v1

@@ 5.3 Qualitative Analysis of Questions
Table 4: The average number of questions asked by different models in an interaction.
Takeaway: Models that balance specificity and question quantity, like the Claude models, achieve greater information gain and superior interaction quality compared to models that ask too few, too many, or templated questions. While Deepseek benefits from asking numerous detailed questions, it risks overwhelming the user. In contrast, Llama underperforms due to its reliance on generic or irrelevant questions.
6 Limitations
Our study benefits from including both open-weight and proprietary models, as well as models from the same family with different parameterizations, enhancing the generalizability of the findings. However, certain design decisions may affect the experiments.
Ambiguity detection is limited to the first three turns, as LLMs struggle to interact meaningfully if they do not engage early. To assess question quality, we measure changes in the latent vector to capture the information gained, assuming equal importance for all new information—though models may prioritize different details in their solution. The resolve rates in the overall problem solving experiment reflect real-life conditions, where incorrect code is unacceptable, regardless of how close the generated patch is to the solution. However, data leakage could enable some models to make correct assumptions in underspecified settings, inflating resolve rates. Additionally, the user proxy may be more interactive than real-world users, as LLMs are tuned to be helpful. We address this by limiting the number of interaction turns and focusing interactions on the task with detailed system prompts.
7 Related Work
Code Generation Benchmarks
In code generation tasks, ambiguous user instructions hinder the evaluation of code suggestions generated by the model. Since the cause of ambiguity is missing details, clarifying questions become neessary (Mu et al., 2023). Interactive, test-driven workflows mitigate this ambiguity by first generating test cases aligned with user expectations, which users validate before code generation (Lahiri et al., 2023). Extensions of this approach employ runtime techniques to generate, mutate, and rank candidate code suggestions and test cases based on user feedback (Fakhoury et al., 2024). Although effective, these workflows can burden users, highlighting the need to minimize intervention to essential cases.
Interactive ML Systems
In task-oriented settings, ambiguity between generated outputs and user expectations remains a challenge. AmbigNLG addresses this by introducing a taxonomy of instruction ambiguities and applying targeted disambiguation based on the identified ambiguity type (Niwa & Iso, 2024). These ambiguities include unclear output lengths, mandatory keywords, and contextual nuances in instructions. NoisyToolBench (Wang et al., 2024a) offers a dataset for evaluating LLM tool use with ambiguous instructions, though it focuses on simpler tasks. Reinforcement learning frameworks like ReHAC balance user interaction by modeling optimal intervention points (Feng et al., 2024), but more effective strategies are needed for complex, multi-step workflows.
LLMs and Ambiguity
The current state-of-the-art LLMs are not inherently trained to handle ambiguity through user interaction (Zhang et al., 2024), but, their instruction tuning enables improved performance with prompt engineering (White et al., 2023). Ambiguity detection has been tackled with uncertainty estimation to measure the utility of seeking clarification (Zhang & Choi, 2023; Park et al., 2024). Meanwhile, the quality of clarifying questions and the resulting performance remain critical to overall success (Rao & Daumé III, 2018; Pyatkin et al., 2023; Kuhn et al., 2023). Despite advances, state-of-the-art techniques such as few-shot prompting and Chain-of-Thought reasoning offer limited relief in ambiguous scenarios (Zhang et al., 2024). Self-disambiguation uses the internal knowledge of a model to reduce query ambiguity (Keluskar et al., 2024; Sterner, 2022; Sumanathilaka et al., 2024). For example, Alignment with Perceived Ambiguity (APA) employs self- disambiguation to quantify perceived ambiguity using information gain, improving the model’s processing of such inputs  (Kim et al., 2024). Although inference-only methods are cost-effective, they are less robust than training-based approaches for handling ambiguity.
8 Conclusion
This work evaluates proprietary and open-weight models in agentic frameworks for handling ambiguity in software engineering. In code generation, to effectively integrate new information into the solution, an agent must detect ambiguity and ask targeted questions. Our key findings are:
  • Given an underspecified input, Claude Sonnet 3.5 and Claude Haiku 3.5 with interaction can achieve 80% of their performance with a well-specified input. In contrast, open-weight models struggle: Deepseek relies on navigational cues to locate relevant files, while Llama 3.1 70B extracts limited information from the user.
  • LLMs do not interact unless explicitly prompted, and their ambiguity detection is highly sensitive to prompt variations. Only Claude Sonnet 3.5 achieves a higher accuracy of 84% in distinguishing between well-specified and underspecified input.
  • Claude Sonnet 3.5, Haiku 3.5, and Deepseek effectively extract new, detailed user information, whereas Llama 3.1 struggles to ask the right questions.
Despite these advances, a gap remains between resolve rates for underspecified vs. fully specified issues. Open-weight models need better interaction strategies to improve resolution, while proprietary models, particularly Claude Haiku 3.5, require stronger prompting to engage interactively. This work establishes the current state-of-the-art in handling ambiguity through interaction, breaking the resolution process into multiple steps.
Impact Statement
This paper examines the ability of AI agents to handle ambiguous and underspecified instructions in software engineering. Our findings highlight key limitations in the resolution of AI-driven tasks in real-world applications, necessitating the development of more interactive, adaptable, and effective systems. By emphasizing AI interactivity, our work contributes to the responsible development of AI technologies, fostering safer and more efficient automation in software engineering.
Appendix A Appendix
A.1 Experimental Design
A.1.1 Full setting
A.1.2 Interaction Setting
A.1.3 Hidden Setting
A.2 Statistical Methods
A.2.1 Wilcoxon Signed-Rank Test
$H_{0}:\tilde{d}\leq 0$
$H_{1}:\tilde{d}>0$
A.3 Ambiguity Detection prompts
A.4 Question Quality Analysis
Cosine $\text{Cosine Distance}(P,Q)=1-\frac{P\cdot Q}{\|P\|\|Q\|}$ (1)
where:
  • $N=1536$ is the dimensionality of the embedding space.
Model Summarized Issue Selected Questions Corresponding Answers
Llama 3.1 70B The InheritDocstrings metaclass is encountering a problem where it fails to handle properties correctly due to its reliance on a function-checking mechanism that does not recognize properties as functions. 1. Are there any existing workarounds or temporary fixes for this problem that I should be aware of? 2. Are there any specific properties or use cases that are causing the issue, or is it a more general problem with the InheritDocstrings metaclass? 1. I don’t have that information. 2. This seems to be a more general problem with how the InheritDocstrings metaclass handles properties, not specific to any properties or use cases.
Deepseek chat The InheritDocstrings metaclass is encountering a problem where it fails to handle properties correctly due to its reliance on a function-checking mechanism that does not recognize properties as functions. 1. Can you provide more details about the InheritDocstrings metaclass? Specifically, what is its purpose, and how is it currently implemented in the repository? 2. Are there any existing tests or examples that demonstrate the issue? If so, can you provide the names of these tests or examples? 1. The InheritDocstrings metaclass is intended to ensure that docstrings are inherited from the parent class when the docstring is not explicitly defined in the child class. It currently uses inspect.isfunction, which doesn’t work with properties since they are not functions. 2. I don’t have that information.
Claude Haiku 3.5 The InheritDocstrings metaclass is encountering a problem where it fails to handle properties correctly due to its reliance on a function-checking mechanism that does not recognize properties as functions. 1. Could you provide more details about where the InheritDocstrings metaclass is located in the repository? 2. When you say the metaclass ”fails to handle properties correctly”, could you elaborate on what specific behavior is incorrect? 1. The InheritDocstrings metaclass is located in the file astropy/utils/misc.py. 2. The incorrect behavior is related to docstring inheritance. Specifically, properties that are redefined in a subclass without a docstring do not inherit the docstring from the base class.
Claude Sonnet 3.5 The issue involves a method that is expected to create a header from a byte string, but it does not accept Python 3 bytes as input. The method works with Python 3’s unicode strings, assuming ASCII encoding, but this behavior is inconsistent with its counterpart method that handles both text and binary file modes. 1. Which specific method in the codebase is causing the issue with byte string input? Please provide the class name and method name. 2. Can you share an example of the byte string input that should work but currently fails? 1. The specific method causing the issue with byte string input is Header.fromstring. 2. I don’t have a specific example of the byte string input, but it should be a byte string that contains the entire header data.
@@ A.4 Question Quality Analysis
Table 5: Examples of questions and their responses across different models chosen to reflect general observed patterns.

### v2

@@ A.6 Question Quality Analysis
Table 7: Examples of questions and their responses across different models reflecting general patterns.
Model Summarized issue Selected questions Answers
Llama 3.1 70B The InheritDocstrings metaclass fails to handle properties due to reliance on function checks that miss properties. 1. Any known workarounds or temporary fixes? 2. Is the issue property-specific or general? 1. I don’t have that information. 2. It seems to be a general problem with how the metaclass handles properties.
Deepseek Chat Same issue with property handling in InheritDocstrings. 1. What is the purpose and current implementation of InheritDocstrings? 2. Are there existing tests that demonstrate the issue? 1. It ensures docstrings are inherited from the parent. Uses inspect.isfunction, which fails for properties. 2. I don’t have that information.
Claude Haiku 3.5 Same issue with improper property handling by InheritDocstrings. 1. Where is the metaclass located in the repo? 2. What is the specific incorrect behavior? 1. It is in astropy/utils/misc.py. 2. Properties redefined without docstrings don’t inherit from the base.
Qwen 3 Coder Same issue with improper property handling by InheritDocstrings. 1. Which classes/modules show the issue? 2. By “properties,” do you mean ‘@property‘? 3. Any tests that already capture the issue? 1. Happens in classes using the metaclass, e.g., in ‘astropy/utils/misc.py‘. 2. Yes, this refers to Python’s ‘@property‘. 3. I don’t have that information.
Claude Sonnet 3.5 A method fails with Python 3 byte input, unlike its counterpart that handles both text and binary modes. 1. Which method has this issue? 2. Example of a failing byte string input? 1. The method is Header.fromstring. 2. No specific example, but it should contain full header data as bytes.
Claude Sonnet 4 Same issue with byte input. 1. Which class/method fails on byte-string input? 2. Example of a byte string that should work but doesn’t? 1. The failing method is ‘Header.fromstring‘. 2. No concrete example, just any full header as a byte string.
A.7 Qwen 3 Coder
Example Qwen 3 thought process in Hidden setting Trajectory message (verbatim): Let me research what the proper RFC standard is. According to RFC 2616 Section 3.3.1: "Recipients of timestamps SHOULD accept all three HTTP-date formats, regardless of the form in which they were sent. For the purposes of HTTP, the preferred format is IMF-fixdate. The other formats are included for compatibility with past and future implementations." And regarding two-digit years: "HTTP/1.1 clients and servers that parse the date value MUST accept all three formats… All HTTP date/time stamps MUST be represented in Greenwich Mean Time (GMT), without exception. For the purposes of HTTP, GMT is exactly equal to UTC (Coordinated Universal Time)… The use of three digit years (e.g. 893) or five digit-plus years (e.g. 17763) is explicitly not allowed." But more importantly, let me check what the actual standard practice is. Looking at common implementations and standards: 1. POSIX/C Standard: Often uses 69–99 $\rightarrow$ 1969–1999, 00–68 $\rightarrow$ 2000–2068 2. RFC 2616 Recommendation: Strongly discourages two-digit years but acknowledges backward compatibility 3. Modern Practice: Many systems use a 50-year sliding window or fixed windows Let me check what the actual issue is referring to. Looking at the existing test and the problem description more carefully, I think the issue is that the current cutoff of 70 is arbitrary and doesn’t follow established standards. Let me check if there are any references or comments in the code that indicate what the intended behavior should be.  Comments: This trajectory message shows Qwen 3 Coder relying heavily on its internal knowledge of RFC standards and POSIX conventions, making specific assumptions about the intended behavior without extensive code exploration. The model directly assumes the cutoff value of 70 is problematic and implements a solution that proves correct. While this could indicate strong pretraining on software engineering standards, this pattern of solving issues through confident assertion rather than exploratory debugging is uncommon in the Hidden setting relative to models that must discover solutions through code analysis.
Qwen 3 Coder and Claude Sonnet 4 Evaluation For both the recent models, we use OpenHands v0.60 and the updated prompts, along with an increased number of steps (up to 100) to account for the increase in capability. For Qwen 3 Coder, we modified the interaction prompt to include a mandatory clarification step, on top of existing interaction instructions. This phase requires the model to output only clarifying questions and wait for responses before proceeding with the problem-solving phases. This modification was necessary because Qwen 3 Coder exhibited a rigid adherence to non-interactive SWE-Bench protocols, often bypassing interaction opportunities even when critical information was missing. The mandatory clarification phase forces the model to engage with the user before attempting implementation. This modification ensures fair comparison in RQ1, which evaluates task success with interaction. Without it, Qwen 3 Coder defaults to non-interactive behavior, invalidating cross-model comparison. RQ2 (detection) and RQ3 (question quality) measure different capabilities and remain unaffected.
A.8 Information Gain
Figure 7: Recovery distance (1 - cosine similarity between full issue and post-interaction knowledge) shows minimal variation across models, failing to capture the extraction efficiency differences revealed by our original metric. This occurs because the full issue contains substantial information (formatting, links, conversational fragments) that is unnecessary for task completion. Models that ask fewer, targeted questions can obtain critical information without recovering irrelevant details, yet are penalized by this metric. In contrast, our extraction-based metric (Figure 5) better captures these differences.